# Removal and Selection: Improving RGB-Infrared Object Detection via Coarse-to-Fine Fusion

### v4

@@ IV-D Comparison with State-of-the-Art Methods
IV-D1 Comparision on the KAIST Dataset
The experiments are conducted on the KAIST under ‘All-dataset’ settings [25, 18]. We compare the proposed RSDet with twelve SOTA multispectral object detection methods on the KAIST dataset.As shown in Table I, We calculate the $\text{MR}^{\text{-2}}$ at IoU values of 0.5 and 0.7, respectively.
When IoU=0.5, according to Table I(a), the RSDet performs best on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of six subsets (‘Near’, ‘Medium’, ‘Far’, ‘None’). And rank second in two subsets (Partial and Heavy). Specifically, the RSDet takes the lead in the subsets of different scales (‘None’, ‘Medium’, ‘Far’), especially in the ‘Far’ subsets, the RSDet surpasses the second-best by about 10%, which is attributed to the mixture of scale-aware experts in the DFS module, Enhancing perceptual ability for targets at different distances. Besides, on the whole test dataset (‘All’), the RSDet surpasses the second-best by about 3%, which indicates the satisfactory accuracy of RSDet. The Log-average Miss Rate over the False Positive Per Image (FPPI) curve of all detectors under IoU=0.5 is shown in Figure 5(a).
@@ IV-D1 Comparision on the KAIST Dataset
When IoU=0.7, according to Table I(b), the RSDet performs best on all the different settings. As IoU increases, the requirement for detection accuracy becomes higher. In this scenario, the superiority of our method becomes even more pronounced. For instance, when IoU is 0.5, all the recent years’ methods MR is 0 under the ’Near’ subset. However, when IoU increases to 0.7, the miss rate of other methods increases more significantly than RSDet, indicating that our proposed method detection results more accurately. The Log-average Miss Rate over the False Positive Per Image (FPPI) curve of all detectors under IoU=0.7 is also shown in Figure 5(b). The above experiment results on the KAIST dataset prove the superiority of our method.
TABLE III: Ablation study on each module result (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in bold.
FLIR LLVIP
RSR DFS $\text{mAP}_{50}$ mAP $\text{mAP}_{50}$ mAP
75.2 37.2 95.0 55.5
76.3 38.4 95.0 57.1
79.8 40.8 95.6 59.5
81.1 41.4 95.8 61.3
IV-D2 Comparision on the FLIR and LLVIP Dataset
The quantitative results of the different methods on the FLIR and LLVIP datasets are shown in Table II. We compare the proposed RSDet with five SOTA multispectral object detection methods and five single-modality object detection methods. From Table II, it can be observed that the experimental results of fusion methods are generally better than those of single-modal methods in the FLIR dataset. However, due to the limited light conditions, the RGB features interfere with the IR features in the LLVIP dataset, resulting in the detection results of the multispectral detector being inferior to the IR modality. However, our method effectively addresses this issue with the well-designed coarse-to-fine fusion strategy and achieves state-of-the-art performance on the both FLIR and LLVIP datasets.
Figure 6: Ablation study on different value $K$ in RSR module on the FLIR and LLVIP dataset.
Figure 7: Qualitative comparison of RSDet with other SOTA methods on the Kaist datasets. From left to right are the Ground Truth on IR and RGB image, MBNet [20], TSFADet [14], CAGTDet [31] and Ours.
Figure 8: Visualization of detection results of our method. The first two rows are from the FLIR dataset and the last row is from the LLVIP dataset. From left to right are the infrared image, the Ground Truth bounding box, and our detection results.
IV-E Ablation Study
IV-E1 Ablation on each component
To verify the effectiveness of the RSR and DFS modules, we ablate each module from RSDet. As shown in Table III, only by adding the RSR module does the performance increase by 1.2% mAP on the FLIR dataset and 1.6% mAP on the LLVIP dataset respectively, and by only adding the DFS module, the metrics show a significant increase on the FLIR and LLVIP dataset, the performance increase by 3.6% mAP on the FLIR dataset and 4.0% mAP on the LLVIP dataset, respectively. Further adding both the RSR and DFS modules, the metrics achieve the best level. The above results indicate the effectiveness of these modules.
IV-E2 Ablation on different K
We further study the impact of different values of $K$ in the RSR module. We conduct the experiments by modifying the value of $K$ ranging from 100 to 400 on the FLIR and LLVIP datasets. As shown in Figure 6, our detector achieves the highest detection performance when $K=320$. If the value of $K$ is increased or decreased, the performance of the detector will decrease, which also proves the negative impact of the redundant spectrum for the detector.
Figure 9: Example of intermediate results of RSR module on FLIR (1,2 rows) and LLVIP (3,4 rows) datasets. “Before” represents the input images, “After” represents the filtered image after the RSR module, and “SNR” is the signal-to-noise ratio. We further visualize the learned filter and the filtered image content in the second column and third column respectively. The last column is the comparison of the background and the foreground region.
Figure 10: tSNE visualization [62] of the modality-specific and shared features. (a) ‘w/o RSR’ represents without RSR module. (b) ‘w RSR’ represents with RSR module.
IV-F Visualization Results
IV-F1 Kaist dataset
We also visualize the object detection results of our RSDet with other state-of-the-art methods on the Kaist testing dataset. From the first two rows of Figure 7, a common issue observed across the MBNet [20], TSFADet [14] and CAGTDet [31] methods is the missed detection of small objects. These results underscore the effectiveness of our DFS module, which can dynamically select different scale features between two modalities. Besides, by cooperating with the RSR module, our method can extract object features more effectively under complex lighting conditions and improve the performance of the detector, which is reflected in the confidence of the bounding box in the last row of Figure 7. The above results verify the effectiveness of our method.
IV-F2 FLIR and LLVIP dataset
In Figure 8, we visualize the object detection results of our method on the FLIR and LLVIP testing dataset. From Figure 8, we can see that our detection results include almost all bounding boxes of the Ground Truth labels with a high degree of consistency. Furthermore, our method can also detect some small objects that are miss-labeled in the Ground Truth, which demonstrates the superiority of our method in multi-scale object detection.
Figure 11: Visualization of DFS module feature fusion results on the FLIR dataset. To facilitate a clearer observation, we overlaid the features onto the original RGB or IR image.
IV-F3 RSR module
We visualize the intermediate results of the RSR module as shown in Figure 9. The second column visualizes the filter used in the frequency domain. Intuitively, both high- and low-frequency information will be filtered. We further transform the filtered information into the spatial domain, visualized in the third column. We can find that the removed information is mainly concentrated in the background region, which illustrates that the RSR module filters out some background redundant noise irrelevant to the detection task. For further objective comparison, we use the signal-to-noise ratio (SNR) to measure the differences, which is calculated using the variance method: $SNR=20\times\log(\frac{\mu}{\sigma^{2}})$. It can be observed that SNR increases after using the RSR module, indicating the effectiveness of the RSR module.
We also conduct tSNE visualizations [62] of shared $C_{\text{sha}}$ and specific features ($C_{\text{I-spe}}$ and $C_{\text{V-spe}}$) on FLIR dataset. As shown in Figure 10, we have observed that without the RSR module, there is still a considerable mix of feature points in the shared and specific features, leading to difficulty in selecting the desired specific features in DFS. After adding the RSR module, we have noted a significant decrease in the number of mixed features, which verifies that removing redundant spectra is beneficial to feature disentanglement and, thus more effective in obtaining fused features in the DFS module.
IV-F4 DFS module
[…]

### v6

@@ IV-D Comparison with State-of-the-Art Methods
IV-D1 Comparision on the KAIST Datas
We compare our proposed RSDet with twelve state-of-the-art methods, on the KAIST dataset. Table IV provides the performance comparisons, we calculate the $\text{MR}^{\text{-2}}$ under IoU thresholds of 0.5 and 0.7. It’s worth noting that the ”All” setting is more challenging, as it includes small (less than 55 pixels) and heavily occluded objects. Our method achieves state-of-the-art results in the ”All” setting, indicating not only strong overall detection performance but also effectiveness at detecting small and heavy occlusion objects.
When IoU = 0.5, according to Table IV(a), RSDet demonstrates outstanding performance, leading in the ”ALL” setting (including all’, day’, and night’) and dominating five of the six subsets (near’, medium’, far’, none’, heavy’). It ranks second only on the partial’ subset. Notably, RSDet exhibits a significant advantage in subsets of different scales (None’, Medium’, Far’), especially in the Far’ subset, where it outperforms the second-best by an impressive margin of approximately 11.72%. We infer that this superior performance can be attributed to the mixture of scale-aware experts in the DFS module, which significantly enhances the model’s ability to perceive objects at different distances. Moreover, across the entire test dataset (‘all’), RSDet surpasses the second-best 4.17%, underscoring its overall superiority.
@@ IV-D1 Comparision on the KAIST Datas
When IoU=0.7, RSDet also performs best across all subsets, as shown in Table IV(b). As the IoU increases, the requirement for detection accuracy becomes more stringent. In this scenario, the superiority of our method becomes more apparent. For example, all the recent methods have an $\text{MR}^{\text{-2}}$ of 0 under the ’near’ subset before the IoU increases to 0.7. However, These methods show an increase in Miss Rate ranging from a maximum of 25.19% to a minimum of 16.98%, whereas our method only rises by 11.81%, the $\text{MR}^{\text{-2}}$ of RSDet is significantly better than other methods, indicating that our proposed method produces more accurate detection results.
To conduct a more comprehensive analysis, we illustrate the Log-average Miss Rate over the False Positive Per Image (FPPI) curve in Figure 10, which highlights the superiority of our method, primarily reflected in the following points: 1) Significant Miss Rate Advantage: RSDet achieves the lowest miss rate among all methods, indicating a stronger ability to identify objects accurately, even under more strict conditions. 2) Excellent False Positive Control: RSDet maintains a notably lower MR even at lower FPPI, demonstrating its effectiveness in reducing false detection while maintaining high detection precision. This enhances the reliability of the detection system. 3) Smooth Curve Performance: The MR-FPPI curve of RSDet is smoother as FPPI increases. This suggests that RSDet offers stable performance across various conditions, making it adaptable to diverse scenarios.
Figure 10: The MR-FPPI curves comparisons with the state-of-the-art methods on the KAIST dataset under the ‘All’ settings.
IV-D2 Comparision on the FLIR Dataset
The quantitative results of the different methods on the FLIR datasets are shown in Table V. We compare the proposed RSDet with twelve SOTA RGB-IR object detection methods on the FLIR dataset. From Table V, it can be seen that the fusion methods generally outperform the single-modal methods on the FLIR dataset. Our RSDet method also stands out, achieving 83.9% $\text{mAP}_{50}$ and 43.8% mAP, significantly surpassing the second-best RGB-IR methods by 4.7% and 2.5%, respectively, demonstrating the superior ability of our method on the FLIR Dataset.
IV-D3 Comparision on the LLVIP Dataset
The quantitative results of the different methods on the LLVIP datasets are shown in Table VI. Due to the LLVIP dataset having only recently been made public, there are relatively few published methods that have conducted experiments on it. Thus, we compare with four SOTA RGB-IR methods which have been compared on FLIR comparison experiments, and the five single-modality methods. In the LLVIP dataset, the poor light conditions result in the RGB features interfering with the IR features during the multimodal feature fusion process, leading to the detection results of the RGB-IR detection method always being outperformed by the single IR modality method. Our method effectively addresses this issue and achieves 95.8% $\text{mAP}_{50}$ and 61.3% mAP, surpassing the second-best RGB-IR methods by 1.2% and 2.1%, respectively. The adequate comparison experiments verify the effectiveness of our coarse-to-fine fusion strategy and achieve state-of-the-art performance on the KAIST, FLIR, and LLVIP datasets.